The chemical equation below shows the complete combustion of propane , C3H8.
C3H8 (g) + 5O2 (g) --> 4H2O(g) + 3CO2 (g)

What is the mass of oxygen required or the complete combustion of 200cm3 of propane measured at r.t.p.?

A. 0.650 g
B. 0.923 g
C. 1.14 g
D. 1.33 g
1
Expert's answer
2015-11-25T09:24:49-0500
Solution
V(C3H8)stp = V(C3H8)stp*273/298 = 0.183
n(C3H8) = V/Vm = 0.183/22.4 = 0.0082 (mol)
n(O2) = n(C3H8) * 5 = 0.041 (mol)
m = n*M = 0.041 * 32 = 1.33 (g)
Answer
D